How do I use logarithms to solve
10(1.07^5x+1)=12

Answer by lwsshak3(11628)

How do I use logarithms to solve
10(1.07^(5x+1)=12
(1.07^(5x+1)=12/10=1.2
take log of both sides
(5x+1)log(1.07)=log(1.2)
5x+1=log1.2/log1.07=2.6947
5x=2.6947-1=1.6947
x=1.6947/5=.3389
Check:
10(1.07^(5x+1)
10(1.07^(2.6947)=10*1.2=12